OpenStudy (stuck-help):

METAL square root 4g^2=25

OpenStudy (owlcoffee):

\[\sqrt{4g^2}=25\] in order to solve for "g" we can choose many ways to solve this, I'll give you two of them so you can see that no matter the method used, the result is the same, it all depends on you knowledge in mathematics. For the first, we will square both sides: \[(\sqrt{4g^2})^{2}=(25)^2\] we can cancel that square root with the squared expression on the left side, and 25^2 is just 625: \[4g^2=625\] And now, let's solve for g: \[g=\sqrt{\frac{ 625 }{ 4 }}\] Well, that's already a number so I have pretty much solved for "g" there. The second is using the radical properties: \[\sqrt{u}\sqrt{p}=\sqrt{up}\] That just means that if I have two radicals multiplying each other and the numbers inside are different, we can just take the radical of their product. This also works backwards: \[\sqrt{4g^2}=25\] \[\sqrt{4}\sqrt{g^2}=25\] and simplifying a little: \[2g=25\] \[g=\frac{ 25 }{ 2 }\] Does not look like it, but this is actually to same solutions for a singular problem, I just used two methods.